Given a macaroni economic model:

C=200+2/3yd

G=500

I=300

X=100

M=20+1/10y

T=15+1/20y

Required: determine equilibrium income, consumption exp, total tax, the multiplier


1
Expert's answer
2021-03-04T15:14:42-0500

In the income-expenditure model, the equilibrium occurs at the level of GDP where aggregate expenditures equal national income (or GDP).

Aggregate expenditure function (AE) equals National Income (Y) 

Y= C+I+G+X-M

"Y=200+2\/3(Y-15-1\/20Y)+300+500+100-20-1\/10Y"

"Y=200+2\/3(19\/20Y-15)+880-1\/10Y"

"Y=1070+32\/60Y"

"Y-32\/60Y=1070"

"28\/60Y=1070"

"Y=2,293" Equilibrium Income.

Consumer Expenditure

"C=200+2\/3(Y-15-1\/20Y)"

"C=200+2\/3(2293-15-(1\/20*2293)"

"C=200+2\/3(2163.35)"

"C=1,642.23"

Total Tax

"T=15+1\/20Y"

"T=15+1\/20*2,293"

"T=129.65"

Multiplier

"\alpha"="1\/(1-c)")

c=0.67

"\alpha"= "1\/(1-0.67)"

"\alpha"= 3.00